why is amount of electricity we need to produce 32 oxygen by electrolysis is the same needed to produce 2 g of hydrogen gas ?
1
Expert's answer
2017-06-06T05:09:34-0400
Solution:
The amount of oxygen is equal to the amount of hydrogen.
n(O_2 )=m/M=32/32=1 mol;
n(H_2 )=m/M=2/2=1 mol.